OpenStudy (anonymous):

simplify: 3^2 · 3^4 A) 3^6 B) 3^2 C) 9^6

OpenStudy (anonymous):

@pooja195

OpenStudy (crabbyoldgamer):

When multiplying a number with the same base to different powers, add the powers. \[3^2 * 3^4 = 3^(2+4) = 3^6\]